# Linear Bandits with Partially Observable Features

### v1

@@ Abstract
We introduce a novel linear bandit problem with partially observable features, resulting in partial reward information and spurious estimates. Without proper address for latent part, regret possibly grows linearly in decision horizon $T$, as their influence on rewards are unknown. To tackle this, we propose a novel analysis to handle the latent features and an algorithm that achieves sublinear regret. The core of our algorithm involves (i) augmenting basis vectors orthogonal to the observed feature space, and (ii) introducing an efficient doubly robust estimator. Our approach achieves a regret bound of $\widetilde{O}(\sqrt{(d+d_{h})T})$, where $d$ is the dimension of observed features, and $d_{h}$ is the unknown dimension of the subspace of the unobserved features. Notably, our algorithm requires no prior knowledge of the unobserved feature space, which may expand as more features become hidden. Numerical experiments confirm that our algorithm outperforms both non-contextual multi-armed bandits and linear bandit algorithms depending solely on observed features.
Keywords: linear bandits, partially observable features, doubly robust estimation
@@ 1 Introduction
We consider a linear bandit problem where the learning agent has access to only a subset of the features, while the reward is determined using the complete set of features, including both observed and unobserved elements  [^1] In this paper, we use the terms “unobserved” and “latent” interchangeably to refer to features that are not visible to the decision-making agent.. Conventional linear bandit problems rely on the assumption that the rewards are linear to only observed features, without accounting for the potential presence of unobserved features. However, in many real-world applications, rewards are often affected by the latent features that are not observable to the agent. For example, in recommendation systems, the true reward — such as user satisfaction or purchase decisions — depends not only on observed features like user demographics or past behaviors but also on latent preferences, such as specific tastes in artists (for streaming services) or brands (in e-commerce). Accurately incorporating these latent features is essential for providing precise recommendations, while ignoring them causes bias or model misspecification errors in every decision-making.
@@
To address the latent features, Park and Faradonbeh (2022), Kim et al. (2023a) and Park and Faradonbeh (2024) rely on the assumption that observed features are linear to the latent features sampled from a specific distribution, e.g., a mean-zero Gaussian. Establishing a regret bound sublinear in the decision horizon without such structural assumptions on the latent features remains a significant challenge and has not been accomplished yet. Key challenges in the bandit problem with partially observable features arise from the complete lack of information on the latent features. Indeed, we do not even know whether an agent observes features partially nor whether we should use the latent features. To address these challenges, we propose a novel linear bandit algorithm that is agnostic to partially observability. Despite the absence of knowledge regarding unobserved features, our algorithm attains a tighter regret bound than both linear bandit algorithms that considers only observed features and MAB algorithms that ignore features entirely. Specifically, our proposed algorithm achieves a $\sqrt{T}$-rate regret bound, without requiring any prior knowledge of the unobserved features, where $T$ is the decision horizon.
@@
The key idea of our proposed algorithm lies in two main components: (i) reconstructing feature vectors to capture the influence of unobserved features on rewards, and (ii) constructing a novel doubly robust estimator to mitigate information loss due to unobservability. For (i), we decompose reward into two additive terms: one projected onto the row space of observed features, and the other onto its orthogonal complement. The former term maximally captures the effect of observed features, while the latter minimizes the impact of unobserved features. We then augment observed features with an orthogonal basis from the complement space. This allows us to reformulate the problem within a conventional linear bandit framework, where the reward function is defined as a dot product of minimally augmented features and an unknown parameter. However, since these augmented features are not identical to the unobserved features, potential estimation error may arise. To mitigate these errors, we leverage (ii) a doubly robust estimator, which is widely used in statistical literature for its robustness to errors due to missing data. These two approaches allow our algorithm to effectively compensate for missing information, improving both estimation accuracy and adaptability to the environment.
Table 1: An overview of regret bound range of our algorithm, RoLF, depending on $d_{h}\in[0,K-d]$, the dimension of the vector space spanned by the rows of the matrix of unobserved features influencing the reward. Note that $\widetilde{O}$ denotes the big-O notation omitting logarithm factors.
Feature space Regret bound
span(observed features) $\supseteq$ span(latent features) $\widetilde{O}(\sqrt{dT})$
span(observed features) $\subseteq$ span(latent features) $\widetilde{O}(\sqrt{KT})$
otherwise $\widetilde{O}(\sqrt{(d+d_{h})T})$
@@
  • We introduce a novel estimation strategy by (i) efficiently augmenting the features that maximally captures the effect of reward projected onto the observed features, while minimizing the impact of unobserved features (Section 4), and (ii) constructing a doubly robust (DR) estimator that mitigates errors from unobserved features. By integrating augmented features with the DR estimator, we guarantee a $t^{-1/2}$ convergence rate on the rewards for all arms in each round $t$ (Theorem 2).
  • We propose an algorithm named Robust to Latent Features (RoLF) for general linear bandit framework with latent features (Algorithm 1). The algorithm achieves a regret bound of $\widetilde{O}(\sqrt{(d+d_{h})T})$ (Theorem 3), where $d_{h}$ is the dimension of the subspace formed by projecting the reward from unobserved features onto the orthogonal complement of the row space of observed features (Section 4.2) and $\widetilde{O}(\cdot)$ is the Big-O notation omitting logarithmic factors. RoLF requires no prior knowledge or modeling of unobserved features yet achieves a sharper regret bound than both linear bandit algorithms that consider only observed features (Li et al., 2010; Abbasi-yadkori et al., 2011; Agrawal and Goyal, 2013; Kim and Paik, 2019) and MAB algorithms (Auer et al., 2002), as summarized in Table 1.
  • Our experiments confirm that our algorithm consistently outperforms MAB and linear bandit algorithms that solely depend on observed features, validating both its practicality and theoretical guarantees.
@@ 2 Related Works
In bandit problems, the learning agent learns only from the outcomes of chosen actions, leaving unchosen alternatives unknown (Robbins, 1952). This constraint requires a balance between exploring new actions and exploiting actions learned to be good, known as the exploration-exploitation tradeoff. Efficiently managing this tradeoff is crucial for guiding the agent towards the optimal policy. To address this, algorithms based on optimism in the face of uncertainty (OFU) (Lai and Robbins, 1985) are widely used and studied in linear bandits (Abe and Long, 1999; Auer, 2002; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2010). Notable examples include LinUCB (Li et al., 2010; Chu et al., 2011) and OFUL (Abbasi-yadkori et al., 2011), known for their practicality and performance guarantees. However, existing approaches differ from ours in two key aspects: (i) they assume that the learning agent can observe the entire feature vector related to the reward, and (ii) their algorithms have regret that scales linearly with the dimension of the observed feature vector, i.e., $\widetilde{O}(d\sqrt{T})$.
In contrast, we develop an algorithm that achieves a sublinear regret bound by employing the doubly robust (DR) technique, thereby avoiding the linear dependence on the dimension of the feature vectors. The DR estimation in the framework of linear contextual bandits is first introduced by Kim and Paik (2019) and Dimakopoulou et al. (2019), and subsequent studies improve the regret bound in this problem setting by a factor of $\sqrt{d}$ (Kim et al., 2021; 2023b). A recent application (Kim et al., 2023c) achieves a regret bound of order $O(\sqrt{dT\log T})$ under IID features over rounds. However, the extension to non-stochastic or non-IID features remains an open question. To address this issue, we develop a novel analysis that applies the DR estimation to non-stochastic features, achieving a regret bound sublinear with respect to the dimension of the augmented feature vectors. Furthermore, we extend DR estimation to handle sparse parameters, thereby further improving the regret bound to be sublinear with respect to the reduced dimension.
Our problem is more general and challenging than misspecified linear bandits, where the assumed reward model fails to accurately reflect the true reward, such as when the true reward function is non-linear (Lattimore and Szepesvári, 2020), or a deviation term is added to the reward model (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022). While our work assumes that the misspecified (or inaccessible) portion of the reward is linearly related to certain unobserved features, misspecified linear bandit problems can be reformulated as a special case of our framework. While the regret bounds in Lattimore and Szepesvári (2020), Bogunovic et al. (2021) and He et al. (2022) incorporate the sum of misspecification errors that may accumulate over the decision horizon, our work establishes a regret bound that is sublinear in the decision horizon $T$, not affected by misspecification errors. Ghosh et al. (2017) proposed a hypothesis test to decide between using linear bandits or MAB, demonstrating an $O(K\sqrt{T}\log T)$ regret bound when the total misspecification error exceeds $\Omega(d\sqrt{T})$. In contrast, our algorithm achieves an $O(\sqrt{(d+d_{h})T\log T})$ regret bound without requiring hypothesis tests for misspecification or partial observability.
Lastly, our problem appears similar to the bandits with partially observable features, as studied by Park and Faradonbeh (2022) and Zeng et al. (2024). Both works assume that observed features are related to latent features through a known linear mapping, with the latent features sampled from a certain Gaussian distribution. Notably, the latter work further assumes that the latent features evolves following a specific linear dynamics model. In addition, both approaches aim to recover the latent features: the former introduces a known decoder mapping from observed features to their corresponding latent features, while the latter estimates them using a Kalman filter. In contrast, our approach imposes no structural assumptions on either the observed or latent features, making our problem more general and fundamentally more challenging than those addressed in these works. Furthermore, our work does not attempt to recover any information related to latent features. Instead, our novel approach demonstrates that the best action can still be identified and chosen, even with an unobserved portion of the reward, by solely exploiting observed features.
@@ 3.2 Problem Formulation
$\mathbf{z}_{a}:={\left[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}x_{a}^{(1)}},\cdots,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}x_{a}^{(d)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}u_{a}^{(1)}},\cdots,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}u_{a}^{(d_{u})}}\right]}^{\top},$ (1)
For clarity, we highlight the observed components in blue, while the unobserved components in red. Note that the dimensions of the latent feature vector, $d_{u}=d_{z}-d$, and the true feature vector, $d_{z}$, are both unknown to the agent. Consequently, the agent is unaware of whether the features are partially observed, which introduces significant challenges in selecting appropriate strategies.
It is worth noting that the setting with fixed observed features [^2] This assumption is standard in linear bandits with model misspecification (Ghosh et al., 2017; Lattimore et al., 2020), which is a special case of our partially observable feature setting. includes linear bandits with misspecification error (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022) as special cases. In Appendix B, we present a setting with varying observed features and an algorithm that achieves $\sqrt{T}$-rate regret bound. Moreover, if latent features were allowed to change arbitrarily over time, the problem would become non-learnable and thus ill-posed. Consequently, assuming fixed features is both natural and well-justified (see Table 2 for comparisons).
Table 2: Summary of problem settings covered in this paper and the corresponding results. Note that if latent features arbitrarily change over time, the problem itself would become non-learnable, making the problem ill-posed (see Appendix B for details).
@@
The reward associated with each arm is defined as the dot product of its true features $\mathbf{z}_{a}$ and an unknown parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$, given by $y_{a,t}=\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle+\epsilon_{t}=\mathbf{z}_{a}^{\top}\boldsymbol{\theta}_{\star}+\epsilon_{t}$ for all $a\in[K]$. The error term, $\epsilon_{t}$, captures the inherent randomness in the reward, and we adopt a standard assumption commonly used in bandit problems:
@@
Let $\{\mathcal{F}_{t}\}_{t\in[T]}$ denote history at round $t$, represented by a filtration of sigma algebras. The reward noise $\epsilon_{t}$ is assumed to be a $\sigma$-sub-Gaussian random variable conditioned on $\mathcal{F}_{t}$. Formally,
@@
$\mathbb{E}[{\exp(\lambda\epsilon_{t})}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]\leq\exp\left(\lambda^{2}\sigma^{2}\over 2\right),$
for all $\lambda\in\mathbb{R}$.
@@
It follows that $\mathbb{E}[{\epsilon_{t}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]=0$, and $\mathbb{E}[{y_{a,t}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]=\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle$ under this assumption. For brevity, we use $\mathbb{E}_{t-1}[\cdot]$ to denote $\mathbb{E}[{\cdot}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]$ henceforth. Given that $\epsilon_{t}$ is sampled after each action is observed, $\epsilon_{t}$ is $\mathcal{F}_{t}$-measurable. To eliminate issues of scale for analysis, we assume that the expected reward $|\langle\mathbf{z}_{a},\boldsymbol{\theta}_{\star}\rangle|\leq 1$ for all $a\in[K]$.
@@
Let $a_{\star}:=\mathop{\mathrm{argmax}}_{a\in[K]}\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle$ denote the optimal action, considering both observed and latent features. The theoretical performance of our algorithm is evaluated through cumulative regret, which measures the total expected difference between the reward of the optimal action and the reward of the action selected in each round. Formally,
@@
$\begin{split}\mathrm{Reg}(T)&=\mathbb{E}\left[\sum_{t=1}^{T}\langle{{\mathbf{z}_{\star}}-{\mathbf{z}_{t}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle\right]\end{split}.$ (2)
@@
Considering the composition of $\mathbf{z}_{a}$ defined in Eq. 1, we can decompose the parameter as $\boldsymbol{\theta}_{\star}=[({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\theta}_{\star}^{(o)}})^{\top},({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{\theta}_{\star}^{(u)}})^{\top}]^{\top}$, where $\boldsymbol{\theta}_{\star}^{(o)}\in\mathbb{R}^{d}$ and $\boldsymbol{\theta}_{\star}^{(u)}\in\mathbb{R}^{d_{u}}$ are the parameters for observed and latent features, respectively. Adopting this, the reward $y_{a_{t},t}$ can also be decomposed into three terms:
$y_{a_{t},t}=\langle{\mathbf{x}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle+\epsilon_{t}+\langle{\mathbf{u}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(u)}}\rangle,$ (3)
@@
where the last term corresponds to the inaccessible portion of the reward. This reward model is equivalent to that imposed in the linear bandits with misspecification error (Lattimore et al., 2020). While the regret bound in Lattimore et al. (2020) includes misspecification error that grows linearly in decision horizon, our proposed method (Section 4) addresses this misspecification error and achieves a regret bound that is sublinear in the decision horizon.
@@
Before presenting our method and algorithm, we first establish a lower bound for the regret incurred by algorithms that disregard the unobserved portion of rewards. Specifically, the following theorem establishes a lower bound for two algorithms relying solely on observed features: OFUL (Abbasi-yadkori et al., 2011) and LinTS (Agrawal and Goyal, 2013).
@@
There exists a problem instance where the expected regret of both OFUL (LinUCB) and LinTS grows linearly in $T$.
@@
Consider a linear bandit problem with action set $\mathcal{A}:=\{1,2\}$, where true feature set is defined as $\mathcal{Z}:=\{[1,3]^{\top},[2,19/4]^{\top}\}\subset\mathbb{R}^{2}$ and the arm 2 is the optimal action. The latent portion of each reward is “large”, in the sense that there exists a constant $C>0$ such that $|\langle{\mathbf{u}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(u)}}\rangle|=|\mathbb{E}[y_{a_{t},t}]-\langle{\mathbf{x}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle|>C$ (following Eq. 3). Under this setup, the estimator associated with observed features is not consistent and the OFUL and LinTS selects the suboptimal arm with probability $\Theta(1)$. Consequently, the regret grows linearly in $T$ for both cases. ∎
Theorem 1 implies that neglecting the latent portion of the reward in decision-making could result in a failure in the learning process of the agent. The comprehensive proof is deferred to Section C.1. While Theorem 1 focuses on OFUL and LinTS, which are known to achieve the most efficient regret bounds for UCB and Thompson Sampling-based policies, we prove an algorithm-agnostic lower bound using different analysis (see Appendix D for details).

### v2

@@ Abstract
We study the linear bandit problem that accounts for partially observable features. Without proper handling, unobserved features can lead to linear regret in the decision horizon $T$, as their influence on rewards is unknown. To tackle this challenge, we propose a novel theoretical framework and an algorithm with sublinear regret guarantees. The core of our algorithm consists of: (i) feature augmentation, by appending basis vectors that are orthogonal to the row space of the observed features; and (ii) the introduction of a doubly robust estimator. Our approach achieves a regret bound of $\widetilde{O}(\sqrt{(d+d_{h})T})$, where $d$ denotes the dimension of the observed features, and $d_{h}$ represents the number of nonzero coefficients in the parameter associated with the reward component projected onto the subspace orthogonal to the row space spanned by the observed features. Notably, our algorithm requires no prior knowledge of the unobserved feature space, which may expand as more features become hidden. Numerical experiments confirm that our algorithm outperforms both non-contextual multi-armed bandits and linear bandit algorithms depending solely on observed features.
Linear Bandits, Partially Observable Features, Doubly Robust, Machine Learning, ICML
@@ 1 Introduction
We consider a linear bandit problem where the learning agent has access to only a subset of the features, while the reward is determined using the complete set of features, including both observed and unobserved elements. Conventional linear bandit problems rely on the assumption that the rewards are linear to only observed features, without accounting for the potential presence of unobserved features. However, in many real-world applications, rewards are often affected by the unobserved—hence latent—features that are not observable to the agent. For example, in online advertising without personalization—i.e., when serving general users—each advertisement is associated not only with observable features such as its category, format, or display time, but also with unobservable factors such as emotional appeal, creative design quality, and brand familiarity. These latent factors influence users’ click-through rates, yet they are not directly quantifiable. Similarly, in clinical trials, treatment outcomes depend not only on observable features like dosage and formulation, but also on unobserved factors such as manufacturing variability and potential side effect risks. Hence, accurately incorporating latent features is essential for providing precise recommendations.
@@
To address latent features,  Park & Faradonbeh (2022, 2024); Kim et al. (2023a); Zeng et al. (2025) typically assume that the true features follow a specific distribution, such as a Gaussian distribution. Establishing a regret bound sublinear in the decision horizon without such structural assumptions on the latent features remains a significant challenge and has not been accomplished yet. Key challenges in the bandit problem with partially observable features arise from the complete lack of information on the latent features. Indeed, we do not even know whether an agent observes features partially nor whether we should use the latent features. We tackle these challenges by proposing a novel linear bandit algorithm that is agnostic to partially observability. Despite having no knowledge of the unobserved features, our algorithm achieves a tighter regret bound than both linear bandit algorithms that consider only observed features and multi-armed bandit (MAB) algorithms that ignore features entirely. For brevity, we will refer to linear bandit algorithms relying solely on observed features as “OLB algorithms” henceforth. Specifically, our proposed algorithm achieves a $\sqrt{T}$-rate regret bound, without requiring any prior knowledge of the unobserved features, where $T$ is the decision horizon.
@@
The key idea of our proposed algorithm can be summarized as the following two procedures: (i) reconstructing feature vectors via feature augmentation to capture the influence of unobserved features on rewards, and (ii) introducing a novel doubly robust estimator to mitigate information loss due to unobservability. For (i), we decompose the reward into two additive components: one projected onto the row space spanned by the observed features, and the other onto its orthogonal complement. The former term maximally captures the effect of observed features, while the latter minimizes the impact of unobserved features. We then augment observed features with an orthonormal basis from the complementary space, which is orthogonal to the row space of observed features. This formulation enables us to reformulate the problem within a conventional linear bandit framework, where the reward function is defined as a dot product of the augmented features and an unknown parameter, without any additional additive term.
However, since these augmented features are not identical to the unobserved features, potential estimation error may arise due to information loss. To mitigate such errors, we leverage (ii) a doubly robust estimator, which is widely used in the statistical literature for its robustness to errors caused by missing data. These two approaches allow our algorithm to compensate for missing information due to partial observability, improving both estimation accuracy and adaptability to the environment.
@@
  • We introduce a novel estimation strategy by (i) augmenting the features that maximally captures the effect of reward projected onto the observed features, while minimizing the impact of unobserved features (Section 4), and (ii) constructing a doubly robust (DR) estimator that mitigates errors from unobserved features. By integrating augmented features with the DR estimator, we guarantee a $t^{-1/2}$ convergence rate on the rewards for all arms in each round $t$ (Theorem 2).
[…]
  • Our numerical experiments confirm that our algorithm consistently outperforms OLB (Li et al., 2010; Agrawal & Goyal, 2013; Kim & Paik, 2019) and MAB (Lattimore & Szepesvári, 2020) algorithms, validating both its practicality and theoretical guarantees (Section 6).
@@ 2 Related Works
While our setting appears similar to prior works on bandit problems with (i) model misspecification and (ii) partial observability, it differs from both lines of research in several aspects, including the nature of the unobserved features and the strategies used to address the problem.
First, our problem setting is more general and challenging than misspecified linear bandit problems, where the reward model deviates from the true reward due to non-linearity (Lattimore & Szepesvári, 2020) or additive deviation terms (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022). While prior works incorporate cumulative misspecification error into the regret bound, we obtain a regret bound that remains unaffected by such deviations. In particular, Ghosh et al. (2017) employ hypothesis testing for model selection and obtain regret bound of $O(K\sqrt{T}\log T)$ under high misspecification. In contrast, our algorithm achieves $O(\sqrt{(d+d_{h})T\log T})$ regret without such tests, while handling partial observability in a unified framework.
Second, in bandit problems with partially observable features, prior works often rely on structural assumptions. For example, Park & Faradonbeh (2022, 2024); Kim et al. (2023a) assume the true features are drawn from a Gaussian distribution, while Zeng et al. (2025) model them as evolving according to a linear dynamical system with additive Gaussian noise, where observed features are generated via a known linear mapping, also corrupted by Gaussian noise. These methods typically aim to recover the true features—via decoders (Park & Faradonbeh, 2022, 2024), Bayesian oracles (Kim et al., 2023a), or Kalman filtering (Zeng et al., 2025). In contrast, our approach makes no structural assumptions about the relationship between observed and unobserved features and does not attempt to recover the latter. Instead, we mitigate the information loss from partial observability by projecting the latent part of the reward using only the observed features. A more detailed and comprehensive literature review is provided in Appendix A.
@@ 3.2 Problem Formulation
$\mathbf{z}_{a}:={\left[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}x_{a}^{(1)}},\cdots,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}x_{a}^{(d)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}u_{a}^{(1)}},\cdots,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}u_{a}^{(d_{u})}}\right]}^{\top}.$ (1)
For clarity, we highlight the observed components in blue and the unobserved components in red. Note that the dimensions of the latent feature vector, $d_{u}=d_{z}-d$, and the true feature vector, $d_{z}$, are both unknown to the agent. As a result, the agent is unaware of how many features—if any—are unobserved, or even whether partial observability exists at all. This lack of structural information presents a fundamental challenge, as it prevents the agent from explicitly modeling or compensating for the unobserved components when selecting an appropriate learning strategy.
It is worth noting that the setting with fixed observed features [^2] This assumption is standard in linear bandits with model misspecification (Ghosh et al., 2017; Lattimore et al., 2020), which is a special case of our partially observable feature setting. includes linear bandits with misspecification error (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022) as special cases. In Appendix D, we present a setting with varying observed features and an algorithm that achieves $\sqrt{T}$-rate regret bound. Moreover, if latent features were allowed to change arbitrarily over time, the problem would become non-learnable and thus ill-posed. Consequently, assuming fixed features is both natural and well-justified (see Table 1 for comparisons).
Table 1: Summary of problem settings covered in this paper and the corresponding results. Note that if latent features arbitrarily change over time, the problem itself would become non-learnable, making the problem ill-posed (see Appendix D for details).
@@
The reward associated with each arm is defined as the dot product of its true features $\mathbf{z}_{a}$ and an unknown parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$, given by
$y_{a,t}=\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle+\epsilon_{t}\qquad\forall a\in[K].$ (2)
Here, $\epsilon_{t}$, represents a random noise term that captures the inherent stochasticity in the reward generation process. We provide an assumption on $\epsilon_{t}$, following the assumption commonly adopted in bandit problems.
@@
Let $\{\mathcal{F}_{t}\}_{t\in[T]}$ denote history at round $t$, represented by a filtration of sigma algebras. The reward noise $\epsilon_{t}$ is assumed to be a $\sigma$-sub-Gaussian random variable conditioned on $\mathcal{F}_{t}$. Formally, for all $\lambda\in\mathbb{R}$,
@@
$\mathbb{E}[{\exp(\lambda\epsilon_{t})}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]\leq\exp\left(\frac{\lambda^{2}\sigma^{2}}{2}\right).$
@@
It follows that $\mathbb{E}[{\epsilon_{t}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]=0$, and $\mathbb{E}[{y_{a,t}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]=\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle$ under this assumption. For brevity, we use $\mathbb{E}_{t-1}[\cdot]$ to denote $\mathbb{E}[{\cdot}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]$ henceforth. Given that $\epsilon_{t}$ is sampled after each action is observed, $\epsilon_{t}$ is $\mathcal{F}_{t}$-measurable. To eliminate issues of scale for analysis, we assume that the expected reward $|\langle\mathbf{z}_{a},\boldsymbol{\theta}_{\star}\rangle|\leq 1$ for all $a\in[K]$, and we do not assume any bound on the norm of $\mathbf{z}_{a}$ for all $a\in[K]$.
@@
Let $a_{\star}:=\mathop{\mathrm{argmax}}_{a\in[K]}\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle$ denote the optimal action, considering both observed and latent features, and $a_{t}$ denote an action chosen in round $t$. The theoretical performance of our algorithm is evaluated through cumulative regret, which measures the total expected difference between the reward of the optimal action and the reward of the chosen action:
@@
$\mathrm{Reg}(T)=\mathbb{E}\left[\sum_{t=1}^{T}\langle{{\mathbf{z}_{a_{\star}}}-{\mathbf{z}_{a_{t}}}},{\boldsymbol{\theta}_{\star}}\rangle\right].$
@@
Considering the composition of $\mathbf{z}_{a}$ defined in Eq. 1, we can decompose the parameter as $\boldsymbol{\theta}_{\star}=[({\boldsymbol{\theta}_{\star}^{(o)}})^{\top},({\boldsymbol{\theta}_{\star}^{(u)}})^{\top}]^{\top}$, where $\boldsymbol{\theta}_{\star}^{(o)}\in\mathbb{R}^{d}$ and $\boldsymbol{\theta}_{\star}^{(u)}\in\mathbb{R}^{d_{u}}$ are the parameters for observed and latent features, respectively. Adopting this decomposition of $\boldsymbol{\theta}_{\star}$, the reward $y_{a_{t},t}$ defined in Eq. 2 can be decomposed into the following three terms:
$y_{a_{t},t}=\langle{\mathbf{x}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle+\epsilon_{t}+\langle{\mathbf{u}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(u)}}\rangle.$ (3)
@@
The last term of Eq. 3 corresponds to the inaccessible portion of the reward. This reward model is equivalent to that imposed in the linear bandits with misspecification error (Lattimore et al., 2020). While the regret bound in Lattimore et al. (2020) includes misspecification error that grows linearly in decision horizon, our proposed method (Section 4) addresses this misspecification error and achieves a regret bound that is sublinear in the decision horizon.
@@
Before presenting our method and algorithm, we first establish a lower bound for the regret incurred by algorithms that disregard the unobserved portion of rewards. Specifically, the following theorem establishes a lower bound for two OLB algorithms: OFUL (Abbasi-yadkori et al., 2011) and LinTS (Agrawal & Goyal, 2013).
@@
Under partial observability, there exists a problem instance where both OFUL and LinTS suffer from cumulative expected regret that grows linearly with $T$ due to their disregard for the unobserved components.
@@
Consider a linear bandit problem with an action set $\mathcal{A}:=\{1,2\}$, where the true feature set is defined as $\mathcal{Z}:=\{[1,3]^{\top},[2,19/4]^{\top}\}\subset\mathbb{R}^{2}$ and arm 2 is the optimal action. The latent portion of each reward is “large”, in the sense that there exists a constant $C>0$ such that $|\langle{\mathbf{u}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(u)}}\rangle|=|\mathbb{E}_{t-1}[y_{a_{t},t}]-\langle{\mathbf{x}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle|>C$, following the decomposition described in Eq. 3. Under this setup, the estimator based solely on observed features is not consistent, and the OFUL and LinTS select the suboptimal arm with probability $\Theta(1)$. Consequently, the regret grows linearly in $T$ in both cases. ∎
Theorem 1 implies that neglecting the latent portion of the reward in decision-making could result in a failure in the learning process of the agent. The comprehensive proof is deferred to Section E.1. While Theorem 1 focuses on OFUL and LinTS, which are known to achieve the most efficient regret bounds for UCB and Thompson Sampling-based policies within the linear bandit framework, we prove an algorithm-agnostic lower bound using different analysis (see Appendix F for details).